OpenStudy (jamanorthi):

Find all the values X for the equation below (x+1)^2-25/9=0

OpenStudy (anonymous):

start with \[(x1)^2=\frac{25}{9}\] then take the square root of both sides, don't forget the \(\pm\)

OpenStudy (anonymous):

oops i meant \[(x+1)^2=\frac{25}{9}\]

OpenStudy (anonymous):

you know the square root of 25 right? also the square root of 9?

OpenStudy (jamanorthi):

yeah

OpenStudy (jamanorthi):

5 and 3

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you now have two equations to solve : \[x+1=\frac{5}{3}\] and also \[x+1=-\frac{5}{3}\] then you are done